What is the electric field at a distance of 3cm from 4mC of negative charge?
The electric potential energy of an object at point A is known to be 50 J. If it is released from rest at A, it gains 30 J of kinetic energy as it moves to point B. What is its potential energy at B?
How much work does a 12-V battery do in pushing 2mC of charge through a circuit containing one light bulb?
What was your electric potential relative to a metal pipe if a spark jumped 1.3 cm (0.5 in.) through dry air from your finger to the pipe?
1
Expert's answer
2020-01-07T12:46:41-0500

Electric field is given by "Q/4\pi\epsilon r^2"

here Q=4*10-3 C and r=3*10-2 m

Electric field= 9*109 * 4*10-3/9*10-4 = 4*1010 V/m


"\Delta K.E.=-\Delta U"

K.E.A - K.E.B = UB - UA

0-30 = UB - 50

UB = 20 J


Work done = charge * potential difference

Work done = 12*0.002 =0.024 J


The electric field is given by volts/distance: E= V/d . The breakdown voltage of dry air is about 3x106 V/m. So solving for V we get

V=Ed

or V=(3*106 V/m)*(0.013m)=39000 V